OpenStudy (anonymous):

Micah paid $12 for 100 photos from the grocery store. Roland paid $18 for 120 photos from the drug store. Write the ratios of the price paid per number of photos purchased for each store. Do the ratios form a proportion? Explain.

OpenStudy (anonymous):

Is this right? how do I fix it? Write the ratios of the price paid per number of photos purchased for each store. 100/12 = 25/3 120/18 = 20/3 Do the ratios form a proportion? Explain. Yes, the ratios to form a proportion, because both denominators equal out to be the same.

OpenStudy (anonymous):

12 : 100 6:50 3:25 18 : 120 6:40 3:20 that's the ratios in its lowest form. I believe

OpenStudy (anonymous):

Okay! I still need help with the last part... Is my answer right?

OpenStudy (anonymous):

id say u r correct, in what u r saying.

OpenStudy (anonymous):

My teacher told me it was incorrect... I don't know how to fix it... I know the first put is right but the second isn't

OpenStudy (anonymous):

not sure then, id assume that 3:20 3:25 is correct. justifying that will take sum research, type the start of your question into google, see what comes up, if u r happy to wait 4 someone smarter than me, then wait here.

OpenStudy (anonymous):

I figured it out! My answer was incorrect, the ratios do not form a proportion. But thank you for all your help!